OpenStudy (anonymous):

rewrite 7(40=2) as the sum of two products

jimthompson5910 (jim_thompson5910):

Hint: a(b+c) = a*b + a*c

jimthompson5910 (jim_thompson5910):

Ex: a = 3, b = 4 and c = 5 a(b+c) = a*b + a*c 3(4+5) = 3*4 + 3*5

OpenStudy (anonymous):

i meant 7(40+2)

jimthompson5910 (jim_thompson5910):

i figured that, does my example help?

OpenStudy (anonymous):

kind of but im still a little confused

jimthompson5910 (jim_thompson5910):

You take the outside number (in the example, it's 3) and you multiply it by EVERY number inside. Then you add the products

OpenStudy (anonymous):

ohh now i get it. thanks for the help.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

i got another question

jimthompson5910 (jim_thompson5910):

go for it

OpenStudy (anonymous):

what is the multiplicative inverse of -1/9

jimthompson5910 (jim_thompson5910):

What number multiplies with what you have there will result in an answer of 1?

jimthompson5910 (jim_thompson5910):

(Some number) * (-1/9) = 1 What is the "some number" ?

OpenStudy (anonymous):

isn't 9?

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

or - 1

jimthompson5910 (jim_thompson5910):

9 times -1/9 = -1 so you're close

jimthompson5910 (jim_thompson5910):

it's not -1

OpenStudy (anonymous):

well you're a big help my teacher wont even help me

jimthompson5910 (jim_thompson5910):

that's just messed up...

jimthompson5910 (jim_thompson5910):

you want some number to multiply with -1/9 to give you positive 1...not negative 1 So keep looking

OpenStudy (anonymous):

it s just 9 right?

OpenStudy (anonymous):

not negative

jimthompson5910 (jim_thompson5910):

no, it has to be negative since the two negatives will cancel

jimthompson5910 (jim_thompson5910):

-9 times -1/9 = +1

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

7[6+5+(-6)] = 7[6+(-6)+5] should that be distributive or associative property or communtive or identity property?

jimthompson5910 (jim_thompson5910):

what change is occurring